At a location above the Earth's atomosphere, the maximum electric field of the radiation from the Sun is 5.8 x 103 V/m. Calculate the intensity of this radiation, in units of Watts/m2.
Answer #1

Average intensity (Iavg)of an electro magnetic radiation in

I=\frac{c\varepsilon oE^{2}}{2}

c=3×108m/s

E=5800V/m

\varepsilon o=8.85*10^{-12}C/N.m^{2}

Substituting values in Iavg equation,we get

I=44657.1W/m2--------(ANSWER)